### v2

# Quantum Generative Diffusion Model: A Fully Quantum-Mechanical Model for Generating Quantum State Ensemble

Abstract
Classical diffusion models have shown superior generative results and have been applied to many problems. Exploring these models in the quantum domain can advance the field of quantum generative learning. In this paper, we introduce the Quantum Generative Diffusion Model (QGDM), a simple and elegant quantum counterpart of classical diffusion models.
The core idea of QGDM is that any target quantum state can be transformed into a completely mixed state, which has the highest entropy and maximum uncertainty about the system, through a non-unitary forward process. Subsequently, a trainable backward process can be used to recover the target state from the completely mixed state. The design requirements for QGDM’s backward process include ensuring non-unitarity while maintaining a low number of parameters. To achieve this, we introduce partial trace operations in the backward process to enforce non-unitary. Additionally, we control the number of trainable parameters by using a parameter-sharing strategy and incorporating temporal information as an input in the backward process. Furthermore, we introduce a resource-efficient version of QGDM, which reduces the number of auxiliary qubits while preserving impressive generative capabilities.
Our proposed models exhibit better convergence performance than Quantum Generative Adversarial Networks (QGANs) because our models optimize a convex distance function using gradient descent. Comparative results with QGANs demonstrate the effectiveness of our models in generating both pure and mixed quantum states. Notably, our models achieve 53.03% higher fidelity in mixed-state generation tasks compared to QGANs. These results highlight the potential of the proposed models to tackle challenging quantum generation tasks.
Index Terms:
Variational quantum algorithms, Quantum machine learning, Quantum generative models, Denoising diffusion probabilistic models.
I Introduction
Figure 1: Visualizing the core idea of the proposed models in the Bloch sphere.
Generative models have attracted widespread attention and achieved significant success in academic and industrial fields [1, 2]. Notably, in the past two years, large models such as ChatGPT and Stable Diffusion have shown remarkable performance and creativity in the fields of human-machine dialogue and image generation [3, 4]. These models have made technological breakthroughs and significantly impacted daily life [5]. The core goal of generative models is to learn the underlying probability distributions from a given dataset. Once trained, these models can generate new data that follow the distribution, demonstrating an understanding of the underlying patterns in the data.
Quantum generative models [6], as the quantum counterparts of classical generative models in the field of quantum machine learning [7, 8, 9, 10, 11], have sparked great interest in the academic community. Quantum generative models merge the advantages of quantum computing and classical generative models, offering new possibilities for data generation. There have been several preliminary explorations of various quantum generative models, including the Quantum Generative Adversarial Network (QGAN) [12, 13, 14], the Quantum Circuit Born Machine (QCBM) [15, 16], the Quantum Variational Autoencoder (QVAE) [17], and the Quantum Boltzmann Machine (QBM) [18]. Among them, QGAN, the quantum counterpart of the classical Generative Adversarial Network (GAN) [19], has received considerable attention for its remarkable applications, including, but not limited to, generating quantum state [13, 20, 21], image generation [22, 23, 24], and fitting classical discrete distributions [14, 25, 26]. QGAN provides a pathway for researchers to more effectively simulate and explore complex quantum systems, especially in generating and processing data related to quantum states.
I-A Motivations
Although QGANs are applied in various fields, they inherit challenges in convergence from classical Generative Adversarial Networks. QGANs may experience model collapse, leading to oscillations in the loss of the generator and discriminator without converging to a global optimum [21]. Furthermore, QGANs face a unique challenge in the quantum domain as they cannot achieve convergence in generating mixed states [27]. Specifically, according to Helstrom’s measurement theory, the analytic optimal output of the generator under a minimax optimization strategy is a pure state, which demonstrates its inability to produce a target mixed state. The aforementioned challenges make the training of QGANs difficult and limit their broader application in critical quantum state generation tasks. These limitations highlight the need for developing more effective alternatives to address these complex tasks of quantum state generation.
To achieve this goal, it is important to explore the quantum analogs of classical generative diffusion models [28, 29, 30, 31]. Classical diffusion models have been proven to have superior generative performance in many fields compared to GAN, and they have more friendly convergence properties than classical GANs [32, 33]. Quantum analogs of diffusion models may offer similar benefits by setting the objective function to minimize a convex distance function between the target quantum state and the generated quantum state. Optimizing a convex function using gradient descent typically results in better convergence properties [34]. Therefore, the development of such models is expected to bypass some challenges faced by QGANs and contribute to the field of quantum generative learning. These considerations motivated us to undertake this work.
I-B Contributions
Inspired by classical diffusion models, this paper introduces a novel quantum generative model known as the Quantum Generative Diffusion Model (QGDM) for quantum state generation tasks. The core idea of QGDM is shown in Fig. 1. Any target quantum state ($\rho_{0}$) can be transformed into a completely mixed state ($\rho_{T}$) through a forward (diffusion) process. Subsequently, a trainable backward (denoising) process can be used to recover the target state from the completely mixed state. Our objectives are to define a timestep-dependent forward process, design an effective quantum state backward process, and benchmark our model in comparison to other quantum generative models in pure and mixed state generation tasks. Our contributions are summarized as follows:
  1. 1.
    We propose an elegant quantum version of the diffusion model called the Quantum Generative Diffusion Model (QGDM). The core idea of QGDM is that any target quantum state can be diffused into a completely mixed state by using a forward process. Then, using a learnable backward process, the original target quantum state can be recovered from the completely mixed state. The forward process employs a non-unitary depolarization channel model. To reverse the diffusion process, we introduce a partial trace operation during denoising, which traces out information from the subsystem and makes the backward process non-unitary. Additionally, we use a parameter-sharing strategy and include temporal information as an inputs, allowing the backward processes at all timesteps to share the same parameters.
  2. 2.
    We introduce a resource-efficient version of QGDM (RE-QGDM) to reduce the number of auxiliary qubits in the backward process, thereby enhancing the model’s efficacy in addressing more complex and large-scale problems.
  3. 3.
    Numerical simulation results demonstrate that the proposed models outperform QGANs in generating both mixed states and pure states. Notably, in generating mixed states, QGDM achieves 22.17% higher fidelity compared to QGANs for tasks involving 1 to 4 qubits, while RE-QGDM achieves 64.66% higher fidelity compared to QGANs for tasks involving 1 to 8 qubits.
The remainder of this article is structured as follows. Initially, in Section II, we delve into the existing literature and the foundational work relevant to quantum generative models. Subsequently, in Section III, we present an in-depth exposition of our innovative Quantum Generative Diffusion Model (QGDM) framework. Further, Section IV is dedicated to introducing a Resource-Efficient iteration of the QGDM (RE-QGDM). Section V analyzes the rationale behind the design of the backward process and presents simulation results to support these insights. These analyses and simulation outcomes are poised to aid future research in similar models. The numerical simulations setup and comparative results with other quantum generative models are shown in Section VI. Finally, Section VIII concludes this article. The notations are summarized in Table I.
TABLE I: Notations.
Notation Description
$t$ Timestep
$T$ Total Timesteps of the Forward Process
$|\psi\rangle$ Quantum State
$\rho$ Density Matrix of A Quantum System
$\rho_{0}$ Target Quantum State
$\rho_{t}$ System’s Density Matrix at Timestep $t$
$\rho_{T}$ Completely Mixed State
$\tau_{t}$ Timestep Embedding Quantum State
$\mathbb{I}$ Identity Matrix
$\alpha_{t}$ Noise Parameter Dependent on $t$
$\bar{\alpha}_{t}$ Accumulated Noise Parameter $\bar{\alpha}_{t}=\prod_{i}^{t}{\alpha_{i}}$
$\mathcal{E}(\cdot,\cdot)$ Forward Process Function
$f_{\boldsymbol{\Theta}}(\cdot,\cdot)$ Backward Process Function Parameterized by $\boldsymbol{\Theta}$
$\mathcal{T}(\boldsymbol{\omega},\cdot)$ Timestep Embedding Circuit Parameterized by $\boldsymbol{\omega}$
$U(\boldsymbol{\theta})$ Denoising Circuit Parameterized by $\boldsymbol{\theta}$
$\text{tr}(\cdot)$ Trace Operation
$\text{tr}_{\textit{B}}(\cdot)$ Partial Trace Operation over Subsystem B
$F(\cdot,\cdot)$ Quantum Fidelity Function
$N$ Number of Qubits in the Target State
$N_{\text{time}}$ Number of Qubits in the Timestep Embedding Circuit
$d$ Dimension of Hilbert Space
$\lambda$ Hyperparameter in Loss Function
$A^{\dagger}$ Hermitian Conjugate of the Operator $A$
$\mathcal{L}_{t}$ Loss Function at Timestep $t$
$\mathcal{L}$ Total Loss Function
II Related Works
II-A Preliminary
We first introduce the necessary quantum information backgrounds. For more details, please refer to Ref. [35].
II-A1 Pure Quantum State
In quantum computing, a state of an $N$-qubit can be represented by a vector in a complex Hilbert space $\mathbb{C}^{d}$ with unit length, denoted as $|\psi\rangle\in\mathbb{C}^{d}$, where the ket notation |$\rangle$| ⟩ denotes a column vector and $d=2^{N}$ is the dimension of Hilbert space. The bra notation $\langle\psi|=|\psi\rangle^{\dagger}$ represents a row vector, with $\dagger$ denoting the conjugate transpose. The evolution of a pure state can be represented by applying a quantum circuit or a quantum gate: $|\psi^{\prime}\rangle=U|\psi\rangle$, where $U$ is a unitary operator satisfying $UU^{\dagger}=\mathbb{I}$, and $\mathbb{I}$ is the identity matrix.
II-A2 Mixed Quantum State
A mixed state of a quantum system is a probabilistic mixture of pure states. It can be represented by a density matrix (or density operator): $\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$, where the ensemble of pure states $\{p_{i},|\psi_{i}\rangle\}$ indicating that the system is in state $|\psi_{i}\rangle$ with a probability of $p_{i}$. The density matrix is particularly useful in contexts where the exact state of the system is not known or when considering subsystems of entangled pairs. The evolution of a mixed state is described by $\rho^{\prime}=U\rho U^{\dagger}$.
II-A3 Tensor Product
The Hilbert space of a combined system, composed of two (or more) different quantum systems, is the tensor product of the Hilbert spaces of the subsystems. Specifically, if the states of the subsystems are $|\psi_{A}\rangle$ and $|\psi_{B}\rangle$ for pure states, or $\rho_{A}$ and $\rho_{B}$ for mixed states, the state of the composite system can be represented as $|\psi_{A}\rangle\otimes|\psi_{B}\rangle$ or $\rho_{A}\otimes\rho_{B}$, respectively. The tensor product operation allows us to describe the complete state of the composite system through the individual states of its components.
II-A4 Partial Trace
Suppose we have two quantum systems, A and B. Their composite system state is represented by the density matrix $\rho_{AB}$. Partial trace operation yields a reduced density matrix for subsystem A, calculated by: $\rho_{A}=\text{tr}_{B}(\rho_{AB})$.
II-B Quantum Generative Models
Quantum generative models, a pivotal sector in quantum machine learning, encompass various models including Quantum Circuit Born Machine (QCBM) [15], Quantum Boltzmann Machine (QBM) [18], Quantum Variational Autoencoder (QVAE) [17], and Quantum Generative Adversarial Network (QGAN) [12, 13].
The Quantum Circuit Born Machine (QCBM) uses quantum circuits to generate a classical bit string from specific probability distributions, a concept first proposed by Benedetti et al. [15]. This model has seen substantial advances [16, 36, 37, 38], finding applications in various fields, including Monte Carlo simulations[39], financial data generation[40, 41], and joint distribution learning[40, 42]. The Quantum Boltzmann Machine (QBM), functioning as a quantum counterpart of the classical Boltzmann Machine [43, 44], uses a quantum device to prepare Boltzmann distributions, by replacing the units in the classical Boltzmann Machine with qubits and substituting the energy function with a parameterized Hamiltonian. QBM was first demonstrated by Amin et al. [18], leading to significant model enhancements [45, 46, 47]. Khoshaman et al. [17] introduced the Quantum Variational Autoencoder (QVAE), using the QBM as a prior in Discrete VAE [48], with subsequent developments including annealer-based [49] and gate-based QVAE [50]. In the realm of QGANs, Lloyd et al. [12] developed a comprehensive taxonomy, classifying them by the quantum or classical nature of their components. This was empirically supported by Dallaire-Demers and Killoran [13], who demonstrated the efficacy of QGAN in generating the desired quantum data. Since then, the field has continuously explored applications in quantum state generation [51, 20, 27, 21], image generation [22, 23, 24], and discrete data generation[14, 25, 26]. For an in-depth review of these quantum generative models, please refer to the comprehensive work by Tian et al. [6].
During our research, several papers [52, 53, 54] also proposed quantum analogs to the classical diffusion model. Ref. [52] introduces several strategies and theoretical tools that could potentially be applied to develop quantum versions of diffusion models. However, the paper lacks specific implementation examples and experimental validation. In contrast, our work introduces a detailed concept and delves into technical details, including comparisons with other quantum generative models. Ref.[53] proposes the QuDDPM algorithm to generate individual pure states from a target distribution. Instead, our work aims to generate the average state of the distribution, which is also an objective considered in other research on quantum generative models [13, 20, 27, 21]. Additionally, the backward process in the QuDDPM algorithm uses a different set of parameters at each timestep. In contrast, we use a parameter-sharing strategy, ensuring that the same set of parameters is used for the backward process at every timestep. This approach keeps the total number of parameters constant, regardless of the number of timesteps, thereby avoiding an increase in parameters as the number of timesteps grows. Ref. [54] uses a classical approach for their diffusion process, while the denoising process is performed using a trainable quantum circuit for the denoising task. However, our method is fully quantum mechanical, encompassing both diffusion and denoising processes.
III Quantum Generative Diffusion Model
III-A General Framework
Figure 2: The framework of the proposed Quantum Generative Diffusion Model (QGDM).
Fig. 2 shows a general framework of the proposed Quantum Generative Diffusion Model (QGDM). QGDM includes a forward process (or diffusion process) $\mathcal{E}(\rho_{t-1},t)$ and a backward process (or denoising process) $f_{\boldsymbol{\Theta}}(\rho_{t},t)$. The diffusion process $\mathcal{E}(\rho_{t-1},t)$ takes an $N$-qubit quantum state $\rho_{t-1}$ and a timestep scalar $t$ as input, producing a quantum state $\rho_{t}$ with added noise. Similar to classical diffusion models, in QGDM, a “timestep” refers to a discrete step in the model’s processes where noise is sequentially added to (forward process) or subtracted from a quantum state (backward process). In this paper, $\mathcal{E}(\rho_{t-1},t)$ is modeled as a depolarizing channel. When $t=0$, $\rho_{0}$ represents the target quantum state. In contrast, at $t=T$, $\rho_{T}$ is a completely mixed state. The denoising process employs a model $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ with trainable parameters $\boldsymbol{\Theta}$, which denoises the noisy input quantum state $\rho_{t}$, producing a cleaner quantum state $\rho_{t-1}$. After applying the function $f_{\boldsymbol{\Theta}}(\cdot,\cdot)$ $T$ times to $\rho_{T}$, the target quantum state $\rho_{0}$ is recovered.
Remark 1: Here we discuss the difference between the notations used in this paper and those used in classical diffusion model papers [29, 32, 33]. In quantum information, the quantum operation formalism $\rho^{\prime}=\mathcal{E}(\rho)$ is a powerful tool for describing the evolution of quantum states [35], where $\mathcal{E}$ represents a quantum operation, and $\rho$ and $\rho^{\prime}$ are the initial and final states, respectively. This formalism is especially useful for describing open quantum systems or the stochastic evolution of quantum states. In this paper, we utilize the quantum operation formalism, denoting forward and backward processes as $\mathcal{E}(\rho_{t-1},t)$ and $f_{\boldsymbol{\Theta}}(\rho_{t},t)$, respectively. These notations do not explicitly include the output term. In contrast, in classical diffusion model papers, forward and reverse distributions are defined as $q(x_{t}|x_{t-1})$ and $p_{\boldsymbol{\theta}}(x_{t-1}|x_{t})$. These notations explicitly include the output term ($i.e.$, $x_{t}$ in $q(x_{t}|x_{t-1})$ and $x_{t-1}$ in $p_{\boldsymbol{\theta}}(x_{t-1}|x_{t})$). The difference between the notations in our paper and those in classical diffusion model papers can be summarized as follows. In the classical case, the evolution process of a sample ($x_{t-1}$ or $x_{t}$) is probabilistic. Therefore, the forward or backward process uses conditional probability notation. In the quantum case, the system’s state is described by a density matrix. The evolution of the density matrix is deterministic, but the density matrix itself represents a probabilistic quantum system. Thus, we use the quantum operation formalism to describe the forward and backward processes. This distinction between quantum and classical systems is emphasized to aid readers who may not be familiar with quantum computation in understanding our notations.
III-B Forward Process
In the natural world, diffusion is governed by non-equilibrium thermodynamics, driving systems toward equilibrium [55]. As a system undergoes evolution, its entropy increases until it reaches maximum entropy at equilibrium, indicating that there is no further evolution. The classical diffusion model draws inspiration from this diffusion mechanism [28]. In quantum information theory, the depolarizing channel is considered as a “worst-case scenario” [56]. Under its influence, a quantum state gradually loses its original information, eventually transforming into a completely mixed state. This state is characterized by the highest entropy, representing the pinnacle of randomness in quantum systems. At this point, our uncertainty about the system has reached its maximum.
An $N$-qubit completely mixed state is mathematically represented as $\mathbb{I}/{2^{N}}=\sum\limits_{i}{\frac{1}{{{2^{N}}}}}\left|i\right\rangle\left\langle i\right|$. This denotes that each computational basis state $\left|i\right\rangle$ occurs with an equal probability of $\frac{1}{2^{N}}$. This uniform probability distribution across all computational bases shows the state to be maximally random, similar to the uniform distribution in classical data. At timestep $t$, the diffusion process of QGDM adds depolarizing noise to an $N$-qubit quantum state $\rho_{t-1}$:
$\rho_{t}$ $=\mathcal{E}\left(\rho_{t-1},t\right)$ (1)
$=(1-\alpha_{t})\mathbb{I}/d+\alpha_{t}\rho_{t-1},$
where $d=2^{N}$ denotes the dimension of Hilbert space and $\alpha_{t}\in[0,1]$ is a timestep-dependent scalar. Eq. (1) indicates that the input quantum state $\rho_{t-1}$ remains unchanged with a probability of $\alpha_{t}$, while with a probability of $1-\alpha_{t}$, it is replaced by the completely mixed state $\mathbb{I}/d$. This indicates the loss of original information in the quantum system. Therefore, the depolarizing channel is non-unitary because information is lost when quantum states pass through it, making the process irreversible. By setting an appropriate sequence of $\{\alpha_{t}\}$, a series of depolarizing channel sequences acting on the quantum state $\rho_{0}$ will eventually yield the output $\rho_{T}\approx\mathbb{I}/d$.
III-B1 Efficient Quantum State Transition to Arbitrary Timestep
Eq. (1) mathematically describes a single-step diffusion process. It is inefficient to repeatedly apply $\mathcal{E}$ to obtain a specific timestep’s quantum state $\rho_{t}$. To enhance the transition efficiency from the initial state $\rho_{0}$ to a specific state $\rho_{t}$, we introduce Theorem 1.
Theorem 1: For any $t\in\{1,\dots,T\}$, the relationship between $\rho_{0}$ and $\rho_{t}$ can be described by:
$\rho_{t}=\left(1-\bar{\alpha}_{t}\right)\mathbb{I}/d+\bar{\alpha}_{t}\rho_{0},$ (2)
where $\bar{\alpha}_{t}=\prod\limits_{i=1}^{t}\alpha_{i}$.
Proof:
$\rho_{t}$ $=(1-\alpha_{t})\mathbb{I}/d+\alpha_{t}\rho_{t-1}$ (3)
$=\left(1-\alpha_{t}\right)\mathbb{I}/d+\alpha_{t}\left[\left(1-\alpha_{t-1}\right)\mathbb{I}/d+\alpha_{t-1}\rho_{t-2}\right]$
$=\left(1-\alpha_{t}\alpha_{t-1}\right)\mathbb{I}/d+\alpha_{t}\alpha_{t-1}\rho_{t-2}$
$=\left(1-\alpha_{t}\alpha_{t-1}\right)\mathbb{I}/d$
$\quad+\alpha_{t}\alpha_{t-1}\left[\left(1-\alpha_{t-2}\right)\mathbb{I}/d+\alpha_{t-2}\rho_{t-3}\right]$
$=\left(1-\alpha_{t}\alpha_{t-1}\alpha_{t-2}\right)\mathbb{I}/d+\alpha_{t}\alpha_{t-1}\alpha_{t-2}\rho_{t-3}$
=$\ldots$= …
$=\left(1-\prod\limits_{i=1}^{t}\alpha_{i}\right)\mathbb{I}/d+\prod\limits_{i=1}^{t}\alpha_{i}\rho_{0}$
$=\left(1-\bar{\alpha}_{t}\right)\mathbb{I}/d+\bar{\alpha}_{t}\rho_{0}.$
Thus, Eq. (2) enables the direct transition from $\rho_{0}$ to a quantum state $\rho_{t}$ in any chosen timestep.
III-B2 Noise Schedule
The noise parameter ${\alpha_{t}}$ determines how fast $\rho_{0}$ approaches $\rho_{T}$. The choice of ${\alpha_{t}}$ must ensure that $\bar{\alpha}_{T}\to 0$ and ${\alpha_{t}}\in\left[{0,1}\right]$ for any $t$. We use the cosine noise schedule from the classical diffusion model paper[31] to select $\alpha_{t}$:
$\alpha_{t}=\frac{\bar{\alpha}_{t}}{\bar{\alpha}_{t-1}},$ (4)
where
$\bar{\alpha}_{t}=\frac{g(t)}{g(0)},\quad g(t)=\cos\left(\frac{t/T+s}{1+s}\cdot\frac{\pi}{2}\right)^{2},$ (5)
with $s$ being a small offset hyperparameter. We note that there are other noise schedules, such as the linear noise schedule [29] and the learned noise schedule [57]. Exploring the impacts of other noise schedules on QGDM’s performance, or designing specific noise schedules for QGDM, is an interesting research direction, which we leave for future work.
III-C Backward Process
The objective of the backward process in QGDM is to restore $\rho_{0}$ from $\rho_{T}$. We should address two considerations in the design of the backward process.
Firstly, since the denoising process is essentially the inverse of the non-unitary diffusion process, it must also be non-unitary to effectively denoise. Secondly, in principle, we need to train $T$ distinct denoising processes to deal with $T$ different denoising demands if we model the denoising process in a form $f_{\boldsymbol{\Theta}_{t}}(\rho_{t})$. To reduce the number of trainable parameters in QGDM, we make these processes share the same parameters by using the timestep $t$ as an input. This integration is crucial for shaping the function $f_{\boldsymbol{\Theta}}(\rho_{t},t)$, as it allows the model to effectively utilize temporal information to accurately estimate $\rho_{t-1}$ while maintaining a low number of parameters.
Figure 3: The backward process framework, $f_{\boldsymbol{\Theta}}(\rho_{t},t)$, is composed of a timestep embedding circuit $\mathcal{T}(\boldsymbol{\omega},t^{\prime})$ and a denoising circuit $U(\boldsymbol{\theta})$, where $\boldsymbol{\Theta}=\{\boldsymbol{\omega},\boldsymbol{\theta}\}$.
To endow the denoising process with the aforementioned two characteristics, we achieve non-unitary properties by tracing out part of the subsystem. Additionally, we use a parameterized circuit to embed temporal information into the quantum system. Our design for the denoising process $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ is illustrated in Fig. 3. $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ comprises two modules: a timestep embedding circuit $\mathcal{T}(\boldsymbol{\omega},t^{\prime})$ parameterized by $\boldsymbol{\omega}$ and $t^{\prime}=t\pi/T$, and a denoising circuit $U(\boldsymbol{\theta})$ parameterized by $\boldsymbol{\theta}$, thus trainable parameters $\boldsymbol{\Theta}=\{\boldsymbol{\omega},\boldsymbol{\theta}\}$. A timestep embedding circuit $\mathcal{T}(\boldsymbol{\omega},t^{\prime})$, operating on an auxiliary register with $N_{\text{time}}$ qubits, is utilized to produce state $\tau_{t}$, which carries temporal information. Subsequently, the composite system $\tau_{t}\otimes\rho_{t}$ is fed into the denoising circuit $U({\boldsymbol{\theta}})$. Afterward, all qubits are regrouped into two registers: register A with $N$ qubits and register B with $N_{\text{time}}$ qubits. We trace out the register B to get the predicted denoised quantum state $\tilde{\rho}_{t-1}$.
III-C1 Timestep Embedding Circuit
A critical consideration when designing the denoising process $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ is how to encode the temporal information $t$ into a quantum state. This is a classical-to-quantum encoding challenge, where methods such as qubit encoding [58, 59, 60] or amplitude encoding [61] could be used to embed $t$ into a quantum system. However, the qubit encoding method only maps the timestep sequence uniformly onto a ring on the Bloch sphere, which lacks expressiveness and flexibility. The amplitude encoding method requires a number of gates that is exponential in the number of qubits [62]. To enhance the expressiveness and flexibility of timestep embedding while maintaining an appropriate gate complexity, in this work, we utilize the quantum embedding method [63] to obtain a timestep embedding state $\tau_{t}$ with $N_{\text{time}}$ qubits:
$\tau_{t}=\mathcal{T}(\boldsymbol{\omega},t^{\prime})\left(\left|{\rm{0}}\right\rangle{\left\langle{\rm{0}}\right|}\right)^{\otimes N_{\text{time}}}\mathcal{T}^{\dagger}(\boldsymbol{\omega},t^{\prime}),$ (6)
where scalar $t^{\prime}=t\pi/T$ is the mapping of timestep $t$ into the range [0, $\pi$], and $\boldsymbol{\omega}$ denotes the trainable parameters in the timestep embedding circuit. The architecture of an timestep embedding circuit with $N_{\text{time}}=4$ qubits is shown in Fig. 4. The circuit includes single-qubit gates $Rx$ and $Ry$, and the two-qubit gate $ZZ(\phi)=\exp{(-i\phi(Z\otimes Z)/2)}$. Our proposed method maps temporal information to appropriate positions on the Bloch sphere (as demonstrated by our numerical simulations) while maintaining linear complexity in the number of quantum gates, which is sufficient for our task requirements.
Figure 4: The architecture of the timestep embedding circuit $\mathcal{T}(\boldsymbol{\omega},t^{\prime})$ with trainable parameter $\boldsymbol{\omega}$, and given parameter $t^{\prime}=t\pi/T$. The circuit in the dashed lines is repeated $L_{\mathcal{T}}$ times. $l\in\{1,\dots,L_{\mathcal{T}}\}$.
III-C2 Denoising Circuit
The denoising circuit $U(\boldsymbol{\theta})$ acts on the composite system $\tau_{t}\otimes\rho_{t}$. Subsequently, all qubits are regrouped into subsystems A and B, where they possess $N$ and $N_{\text{time}}$ qubits, respectively. By tracing out subsystem B, we obtain the predicted output $\tilde{\rho}_{t-1}$:
${{{\tilde{\rho}}_{{\rm{t}}-1}}}$ $=f_{\boldsymbol{\Theta}}(\rho_{t},t)$ (7)
$=\text{tr}_{\text{{B}}}\left(U(\boldsymbol{\theta})\left(\tau_{t}\otimes\rho_{t}\right)U^{\dagger}(\boldsymbol{\theta})\right).$
Using register B to output ${\tilde{\rho}}_{t-1}$ might seem intuitive. However, such a design would result in the model failing. This failure is caused by the slight difference between $\rho_{t}$ and the ground truth $\rho_{t-1}$, which poses a training challenge. Rather than producing the intended output, the optimizer often merely copies its input. Section V provides a detailed examination of this issue, including numerical simulations to demonstrate the concept.
The architecture of a 4-qubit denoising circuit $U(\boldsymbol{\theta})$ is shown in Fig. 5. The circuit includes single-qubit gates $Rz$ and $Rx$, and the two-qubit gate $XX(\phi)=\exp{(-i\phi(X\otimes X)/2)}$. This circuit architecture using $\mathcal{O}(N^{2})$ two-qubit gates may affect the efficiency of our method. However, our focus is to evaluate the effectiveness of our method. To ensure that this evaluation is not affected by the circuit design, we choose a more complex architecture, even though it increases the gate complexity. This decision allows our results to accurately represent the potential of our method. In addition, it may not be necessary for us to maintain a consistent circuit architecture across all timesteps. Adaptable circuits that change architecture at different timesteps can reduce the total number of gates in QGDM. We leave the task of designing efficient circuit architectures for QGDM for future work. An effective approach is to use quantum architecture search algorithms [64, 65, 66, 67, 68, 69].
Figure 5: The architecture of a 4-qubit denoising circuit $U(\boldsymbol{\theta})$, where $\boldsymbol{\theta}$ denotes the trainable parameters. To enhance the learning capability of the circuit, the part enclosed by the dashed lines is repeated $L$ times. $l\in\{1,\dots,L\}$.
III-D Training and Generation
The training objective of QGDM is to maximize the quantum fidelity between the predicted denoised state $\tilde{\rho}_{t-1}$ and the ground truth $\rho_{t-1}$. The loss function is defined as follows:
$\min_{\boldsymbol{\Theta}}\mathcal{L}=\mathcal{L}_{0}+\lambda\mathbb{E}_{t\in\text{Uniform}(\{2,\dots,T\})}\left[\mathcal{L}_{t-1}\right],$ (8)
where $\lambda$ is a hyperparameter used to balance the losses $\mathcal{L}_{0}$ and $\mathbb{E}_{t\in\text{Uniform}(\{2,\dots,T\})}\left[\mathcal{L}_{t-1}\right]$, and we find that setting a small $\lambda$ improves the generative effect of QGDM. $\text{Uniform}(\{2,\dots,T\})$ represents a uniform distribution that gives scalar values from 2 to $T$. $\mathcal{L}_{t-1}$ is calculated by:
${\mathcal{L}_{t-1}}=1-F\left(\rho_{t-1},{\tilde{\rho}}_{t-1}\right),$ (9)
where $F(\cdot,\cdot)$ is the quantum fidelity function [35, 56], used to measure the distance between any two quantum states $\rho$ and $\sigma$:
$F(\rho,\sigma)=\left(\operatorname{tr}\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}\right)^{2}.$ (10)
The QGDM training algorithm is presented in Algorithm 1. In each iteration, we compute the $\mathcal{L}_{0}$ and the expectation $\mathbb{E}_{t\in\text{Uniform}({2,\dots,T})}[\mathcal{L}_{t-1}]$ to update the trainable parameters until the training loss converges or the maximum number of iterations is reached. The expected value is estimated using the Monte Carlo approach.
Once the parameters are trained, we can iteratively obtain the target state $\rho_{0}$ from the completely mixed state $\rho_{T}$ using $\boldsymbol{\omega}$ and $\boldsymbol{\theta}$. The target quantum state $\rho_{0}$ is then generated using Algorithm 2. The algorithm starts with a completely mixed state and uses the trained denoising process to generate the target state step by step.
Algorithm 1 QGDM training algorithm.
1:  Input: Noise schedule parameters $\{\bar{\alpha}_{t}\}_{t=1}^{T}$, batch size BS, weight parameter $\lambda$, number of qubits of the target state $N$, number of qubits of the timestep embedding state $N_{\text{time}}$, total timestep of the diffusion process $T$, maximum iteration number, target quantum state $\rho_{0}$, learning rate $\eta$, completely mixed state $\mathbb{I}/d$.
2:  $\boldsymbol{\theta}\sim\text{Uniform}(0,\pi)$, $\boldsymbol{\omega}\sim\text{Uniform}(0,\pi)$
3:  repeat
4:     $\rho_{1}\leftarrow\left(1-\bar{\alpha}_{\mathrm{1}}\right)\mathbb{I}/d+\bar{\alpha}_{\mathrm{1}}\rho_{0}$
5:     $t^{\prime}\leftarrow\pi/T$
6:     $\tau_{1}\leftarrow\mathcal{T}(\boldsymbol{\omega},t^{\prime})\left(\left|{\rm{0}}\right\rangle{\left\langle{\rm{0}}\right|}\right)^{\otimes N_{\text{time}}}\mathcal{T}^{\dagger}(\boldsymbol{\omega},t^{\prime})$
7:     $\tilde{\rho}_{0}\leftarrow\text{tr}_{\text{{B}}}\left(U(\boldsymbol{\theta})\left(\tau_{1}\otimes\rho_{1}\right)U^{\dagger}(\boldsymbol{\theta})\right)$
8:     $\mathcal{L}_{0}\leftarrow 1-F(\rho_{0},\tilde{\rho}_{0})$
9:     $\mathcal{L}\leftarrow 0$
10:     for $i$ from 1 to BS do
11:        Sample a unique timestep $t$ from the uniform distribution $\text{Uniform}(\{2,\dots,T\})$
12:        $\rho_{t}\leftarrow\left(1-\bar{\alpha}_{\mathrm{t}}\right)\mathbb{I}/d+\bar{\alpha}_{\mathrm{t}}\rho_{0}$
13:        $t^{\prime}\leftarrow t\pi/T$
14:        $\tau_{t}\leftarrow\mathcal{T}(\boldsymbol{\omega},t^{\prime})\left(\left|{\rm{0}}\right\rangle{\left\langle{\rm{0}}\right|}\right)^{\otimes N_{\text{time}}}\mathcal{T}^{\dagger}(\boldsymbol{\omega},t^{\prime})$
15:        $\tilde{\rho}_{t-1}\leftarrow\text{tr}_{\text{{B}}}\left(U(\boldsymbol{\theta})\left(\tau_{t}\otimes\rho_{t}\right)U^{\dagger}(\boldsymbol{\theta})\right)$
16:        $\rho_{t-1}\leftarrow\left(1-\bar{\alpha}_{\mathrm{t-1}}\right)\mathbb{I}/d+\bar{\alpha}_{\mathrm{t-1}}\rho_{0}$
17:        $\mathcal{L}\leftarrow\mathcal{L}+1-F\left(\rho_{t-1},\tilde{\rho}_{t-1}\right)$
18:     end for
19:     $\mathcal{L}\leftarrow\mathcal{L}_{0}+\lambda\frac{1}{\text{BS}}\mathcal{L}$
20:     $\boldsymbol{\omega}\leftarrow\boldsymbol{\omega}-\eta\nabla_{\boldsymbol{\omega}}\mathcal{L}$
21:     $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\nabla_{\boldsymbol{\theta}}\mathcal{L}$
22:  until $\mathcal{L}$ converges or the number of iterations reaches the maximum
23:  Output: Optimal parameters $\boldsymbol{\omega}^{*},\boldsymbol{\theta}^{*}$
Algorithm 2 QGDM generation algorithm.
1:  Input: Number of qubits of the timesetp embedding state $N_{\text{time}}$, total timestep of the diffusion process $T$, the optimal parameters $\boldsymbol{\omega}^{*},\boldsymbol{\theta}^{*}$, completely mixed state $\rho_{T}=\mathbb{I}/d$.
2:  Initialize $\tilde{\rho}$ as completely mixed state: $\tilde{\rho}\leftarrow\rho_{T}$
3:  for $t$ from $T$ to 1 do
4:     $t^{\prime}\leftarrow t\pi/T$
5:     $\tau_{t}\leftarrow\mathcal{T}(\boldsymbol{\omega}^{*},t^{\prime})\left(\left|{\rm{0}}\right\rangle{\left\langle{\rm{0}}\right|}\right)^{\otimes N_{\text{time}}}\mathcal{T}^{\dagger}(\boldsymbol{\omega}^{*},t^{\prime})$
6:     Update the generation state: $\tilde{\rho}\leftarrow\text{tr}_{\text{{B}}}\left(U(\boldsymbol{\theta}^{*})\left(\tau_{t}\otimes\tilde{\rho}\right)U^{\dagger}(\boldsymbol{\theta}^{*})\right)$ using the previous value of $\tilde{\rho}$.
7:  end for
8:  Output: The final generation state $\tilde{\rho}_{0}=\tilde{\rho}$
IV Resource-Efficient Quantum Generative Diffusion Model (RE-QGDM)
The number of qubits required for the QGDM denoising process is $N_{\text{time}}+N$. To reduce the number of auxiliary qubits, we propose a new design for the denoising process $f_{\boldsymbol{\Theta}}(\rho_{t},t)$. We call this prototype as the Resource-Efficient Quantum Generative Diffusion Model (RE-QGDM).
Figure 6: The denoising process design in Resource-Efficient Quantum Generative Diffusion Models (RE-QGDM).
In RE-QGDM, only $N+1$ qubits are needed to generate an $N$-qubit quantum state, which enhances its efficiency in terms of resource utilization. Fig. 6 illustrates the denoising process circuit for a 3-qubit target state using RE-QGDM. The polluted state $\rho_{t}$ is compressed by a parameterized circuit $U_{1}(\boldsymbol{\theta}_{1})$, with $\boldsymbol{\theta}_{1}$ as its trainable parameters, condensing its information into the final qubit. This qubit, containing concentrated information, is then combined with $\tau_{t}$ to form a composite system. Another circuit, $U_{2}(\boldsymbol{\theta}_{2})$, processes this expanded state across $N+1$ qubits. After applying $U_{2}(\boldsymbol{\theta}_{2})$, the system’s last qubit is discarded, leaving the reduced density matrix of the remaining $N$ qubits as the predicted output state ${\tilde{\rho}}_{{\rm{t}}-1}$:
$\tilde{\rho}_{t-1}={\rm{t}}{{\rm{r}}_{4}}\left[{U_{2}\left({{\boldsymbol{\theta}_{2}}}\right)\left({{\tau_{t}}\otimes\rho_{t}^{{}^{\prime}}}\right){U_{2}^{\dagger}}\left({{\boldsymbol{\theta}_{2}}}\right)}\right],$ (11)
where $\tau_{t}$ is given by Eq. (6) and $\rho_{t}^{{}^{\prime}}$ is calculated by tracing out the first two qubits of the output from the $U_{1}$ circuit:
$\rho_{t}^{{}^{\prime}}={\rm{t}}{{\rm{r}}_{12}}\left({U_{1}\left({{\boldsymbol{\theta}_{1}}}\right){\rho_{t}}{U_{1}^{\dagger}}\left({{\boldsymbol{\theta}_{1}}}\right)}\right).$ (12)
Therefore, the trainable parameters for the denoising process $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ of RE-QGDM are $\boldsymbol{\Theta}=\{\boldsymbol{\omega},\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}\}$. Both QGDM and RE-QGDM share the same diffusion process, training, and generation pipeline, but they differ in the denoising process.
V Critical Evaluation of Denoising in Quantum Generative Diffusion Models
In this section, we provide a detailed analysis of QGDM’s denoising process and present numerical simulation results to support our arguments.
A straightforward design for the denoising process $f_{\boldsymbol{\Theta}}(\rho_{t},t)$ involves using a trainable circuit $U(\boldsymbol{\theta})$ that acts on the extended system $\tau_{t}\otimes\rho_{t}$ and then traces out the qubits of $\tau_{t}$, which can be described by the formula:
$\text{tr}_{(\text{register of }\tau_{t})}\left(U(\boldsymbol{\theta})\left(\tau_{t}\otimes\rho_{t}\right)U^{\dagger}(\boldsymbol{\theta})\right)=\tilde{\rho}_{t-1}.$ (13)
However, the noise added to the quantum state in a diffusion step is negligible, resulting in only a trivial difference between $\rho_{t}$ and $\rho_{t-1}$. In the case of tracing out the register of $\tau_{t}$ and aiming for $\tilde{\rho}_{t-1}\approx\rho_{t}$, it can mislead the optimizer, which tends to adjust the circuit $U(\boldsymbol{\theta})$ to produce an output that closely approximates its input, thus minimizing the training loss. However, this adjustment deviates from the intended role of the denoising circuit.
The numerical simulation results illustrating the limitations of this design approach are shown in Fig. 7. Part (a) displays the decreasing trend of the training loss over epochs. Part (b) displays the Hilbert-Schmidt (HS) distance [35, 56] between the input and output states of the denoising circuit $U(\boldsymbol{\theta})$ during training. HS distance, a measure used in quantum information to quantify the distance between two quantum states, is defined as $D_{\text{HS}}(\rho,\sigma)=\sqrt{\text{tr}[(\rho-\sigma)^{2}]}$. Part (c) displays the trend of generation fidelity of QGDM over timestep $t$. The orange solid line represents the fidelity of the generated states with the target state, while the blue dashed line represents the theoretical fidelity curve that the model should achieve. This result shows the model fails to change the input state during the generation process.
Fig. 7 (a) and (b) indicate that with increasing epochs, both training loss and HS distance rapidly decrease. A reduction in HS distance shows that the difference between the input state $\tau_{t}\otimes\rho_{t}$ and the output state $U(\boldsymbol{\theta})(\tau_{t}\otimes\rho_{t})U^{\dagger}(\boldsymbol{\theta})$ becomes minimal, ultimately converging to zero. At this point, tracing over the subsystem containing $\tau_{t}$ would yield a result close to $\rho_{t}$. Fig. 7 (c) shows that QGDM cannot alter the input state during the generation process. The denoising process merely outputs the quantum state fed into it, hence the fidelity between the generated quantum state and the target remains unchanged.
Figure 7: (a) Training loss with respect to the iterations of training. (b) Hilbert-Schmidt (HS) distance between the input and output state of denoising circuit $U(\boldsymbol{\theta})$. (c) After training, the generation fidelity of QGDM with respect to the timestep. The blue dashed line represents the theoretical fidelity that QGDM should achieve, showing the change in the generation process fidelity over timestep $t$. The orange solid line represents the actual generation fidelity.
The comparison between the final generated quantum state $\tilde{\rho}_{0}$ (shown in the first row) and the target quantum state $\rho_{0}$ (displayed in the second row) is illustrated in Fig. 8. Given the complex nature of the density matrix, the figure separately draws the real and imaginary parts, with the former on the left and the latter on the right. This visualization clearly shows that $\tilde{\rho}_{0}$ remains equal to the completely mixed state $\rho_{T}$, indicating that $U(\boldsymbol{\theta})$ essentially reproduces the input state. This highlights the critical need for a well-designed denoising process in QGDM that effectively integrates information between the noisy input state $\rho_{t}$ and the timestep embedding state $\tau_{t}$, especially considering the minor differences in quantum states across adjacent timesteps. The results and discussions in our study will provide valuable insights for the future development of similar quantum machine learning models.
Figure 8: Visualization of the generated density matrix and target density matrix.
Finally, we observe that $U(\boldsymbol{\theta})$ is not necessarily optimized to the identity matrix $\mathbb{I}$ every time, although it seems straightforward to use $\mathbb{I}$ as a simple way to keep input and output states the same. Learning algorithms do not always choose to do so. This is because different ensembles of quantum states can be transformed into each other by $U(\boldsymbol{\theta})$ while maintaining the same density matrix. There may be more than one transform in the hypothesis space that can satisfy this condition. After training, which one of them is chosen depends on the position of $U(\boldsymbol{\theta})$ in the hypothesis space when the parameters $\boldsymbol{\theta}$ are initialized.
VI Numerical Simulations
Our evaluation of the proposed models involved generating a variety of random quantum states, encompassing both pure and mixed states. We conducted a comparative analysis of our models in comparison to QuGAN [13] and EQ-GAN [21]. QuGAN implementation follows the original framework in Pennylane [70], whereas EQ-GAN is built using the official source code [71].
VI-A Setup
We use the Tensorcircuit framework [72] to simulate quantum circuits and the Tensorflow framework [73] for parameter optimization, with the optimizer Adam [74]. During training, a larger learning rate may make it difficult for the model to converge, while setting the learning rate too small makes the model converge too slowly. To better optimize the model, we use a cosine decay method [75] to dynamically adjust the learning rate. The hyperparameters of QGDM and RE-QGDM are shown in Appendix A.
Our numerical simulations covered a number of qubits ranging from 1 to 8. QGDM achieved results up to $N=4$. EQ-GAN achieved $N=7$ in generating mixed states, as simulating the $N=8$ task requires an unattainable amount of computer memory. RE-QGDM degenerates into QGDM at $N=1$. Therefore, in the next two subsections, the results for QGDM at $N>4$, RE-QGDM at $N=1$, and EQ-GAN at $N=8$ in mixed-state generation are excluded. All simulations were repeated 10 times with different random seeds to obtain the statistical characteristics of our results.
VI-A1 Target Pure State Preparation
Fig. 9 illustrates the circuit to obtain a target pure state $|\psi\rangle$. All parameters in the target state generation circuit are uniformly sampled from the range $[0,\pi]$.
Figure 9: The architecture of a 4-qubit circuit to prepare a target pure state. All the parameters $\boldsymbol{\theta}$ are initialized in $[0,\pi]$.
VI-A2 Target Mixed State Preparation
Quantum mixed states are generated through $\rho=\sum_{i}{p_{i}\left|\psi_{i}\right\rangle\left\langle\psi_{i}\right|}$, where $\{\left|\psi_{i}\right\rangle\}$ are randomly produced by the circuit shown in Fig. 9. The probability vector $\{p_{i}\}$ is obtained by uniform sampling from $(0,1]$ and normalized through a softmax operation to ensure $\sum_{i}{p_{i}}=1$ and $p_{i}>0$ for any index $i$. In this paper, we generate a target mixed state by probabilistically mixing two random pure states.
VI-B Pure State Generation
VI-B1 The Impact of $N_{\text{time}}$
We first investigated the impact of $N_{\text{time}}$ in QGDM for generating pure states. The results are shown in Fig. 10. Generation fidelity improves with $N_{\text{time}}$ when $N_{\text{time}}\leq N$, achieving a maximum when $N_{\text{time}}=N$ for all $N$.
Figure 10: The generation fidelity of QGDM in generating target pure states for different qubits with respect to $N_{\text{time}}$. All data represent the average of 10 parallel experiments, and the error bars indicate the standard deviation.
However, when $N_{\text{time}}>N$, except for the case $N=1$, fidelity decreases slightly in other cases. The reason for this may be that there is a redundant qubit that complicates the denoising process. The denoising circuit aims to predict the denoising state $\tilde{\rho}_{t}$ by merging information from $\tau_{t}$ and $\rho_{t}$, and then projecting its prediction onto the first $N$ qubits of the $\tau_{t}$ state. These qubits are denoted as register A in Fig. 3. In cases where $N_{\text{time}}>N$, there is an extra qubit in the state $\tau_{t}$ compared to $\rho_{t}$. This additional qubit may be redundant, complicating the process of merging and output prediction.
When $N_{\text{time}}<N$, register A contains a part of the qubits of $\rho_{t}$. However, the presence of overlapping qubits between register A and the qubits of $\rho_{t}$ may diminish the efficacy of $U(\boldsymbol{\theta})$ in such scenarios since a part of the information from $\rho_{t}$ is already stored in register A. In Section V, we investigate a well-intentioned but faulty denoising design. This design involves transferring all qubits from $\rho_{t}$ to the output register, maximizing the overlap between register A and the qubits of $\rho_{t}$. This extreme case strengthens our hypothesis, illustrating the potential implications.
Based on the results in Fig. 10, subsequent simulations default to $N_{\text{time}}=N$ as this setting yields the best generative results with our proposed method.
VI-B2 Pure States Generation Results
TABLE II: Comparison of pure state generation between QuGAN, EQ-GAN, and our proposed methods.
Models $N=1$ $N=2$ $N=3$ $N=4$ $N=5$ $N=6$ $N=7$ $N=8$
QuGAN [13] 0.999$\pm$5e-5 0.996$\pm$4e-3 0.995$\pm$4e-3 0.993$\pm$3e-3 0.990$\pm$4e-3 0.986$\pm$3e-3 0.978$\pm$6e-3 0.971$\pm$1e-1
EQ-GAN [21] 1.00$\pm$0 0.978$\pm$5e-2 0.947$\pm$4e-2 0.896$\pm$8e-2 0.898$\pm$9e-2 0.872$\pm$9e-2 0.833$\pm$1e-1 0.824$\pm$8e-2
QGDM (Ours) 0.999$\pm$3e-5 0.999$\pm$2e-4 0.999$\pm$2e-3 0.996$\pm$8e-3 / / / /
RE-QGDM (Ours) / 0.994$\pm$8e-3 0.999$\pm$6e-4 0.999$\pm$4e-4 0.997$\pm$4e-3 0.995$\pm$1e-2 0.994$\pm$4e-3 0.990$\pm$1e-2
Figure 11: Pure state generation results. The circles represent outliers, which are data points that significantly deviate from the rest of the data.
In Fig. 11, we compare our proposed QGDM and RE-QGDM with QuGAN [13] and EQ-GAN [21] to assess their fidelity in generating the desired pure state. The results indicate that as the number of qubits in the target state increases, the fidelity of all models decreases.
The QuGAN model maintains high fidelity at lower qubit counts. However, its median fidelity decreases as the number of qubits increases, along with an increase in variance. For instances with $N\geq 7$, QuGAN’s fidelity falls below 0.99 for all 10 parallel experiments, indicating its limitations in handling more complex quantum states. The EQ-GAN model achieves a perfect fidelity of 1.0 when the system has only one qubit. However, its fidelity decreases as the number of qubits grows, displaying the largest variance among the four models. From 3 qubits onward, EQ-GAN’s fidelity consistently stays below 0.99, highlighting its challenges in generating pure states. The fidelity of our proposed methods, QGDM and RE-QGDM, remains mostly above 0.99 across all qubit numbers, with more concentrated data distributions, indicating better scalability. In most scenarios, RE-QGDM surpasses the other three models in terms of median fidelity and demonstrates greater stability at higher qubit counts compared to QuGAN and EQ-GAN. In summary, QGDM excels in smaller quantum systems, while RE-QGDM offers a more scalable solution across a broader range of qubit configurations.
To better evaluate the performance differences among models, we present the generative fidelity values for each model in Table. II. The average fidelity of QuGAN diminishes as the number of qubits increases, starting from 0.999$\pm$5e-5 at $N=1$ to 0.971$\pm$1e-1 at $N=8$. It is worth noting that its fidelity drops below 0.99 for $N\geq 6$, indicating a decline in performance as the quantum system size increases. The fidelity of QuGAN from $N=1$ to $5$ remains above 0.99, showing a downward trend for larger systems, although these values are consistently lower than those achieved by our proposed method. The EQ-GAN model achieves the best initial result, producing fidelities of 1 at $N=1$. However, as the number of qubits increases, EQ-GAN shows the lowest average fidelity among the four models, along with the highest variance. QGDM exhibits stable fidelity, with an average of 0.999 at $N=1,2$, and $3$, and 0.996$\pm$8e-3 at $N=4$. Importantly, the RE-QGDM model consistently demonstrates high fidelity across all numbers of qubits, even achieving 0.99 at $N=8$. This underscores the superior robustness and accuracy of RE-QGDM in pure state generation, highlighting the effectiveness and scalability of our models.
VI-B3 Timestep Embedding State Visualization
Fig. 12 shows selected examples visualizing a single-qubit timestep embedding state sequence $\{\tau_{t}\}_{t=1}^{T}$ after training, with $T=30$. Different colors are used to denote the states in each timestep $t$. Each example demonstrates that $\tau_{t}$ shifts its position on the Bloch sphere surface coherently as $t$ progresses. The density of the states varies as well. Fig. 12 (a) and (b) show a more dispersed distribution on the Bloch sphere surface. Subfigures (c) and (d) illustrate cases where states are more clustered. Subfigures (e) and (f) display an intermediate situation, with the state sequence moderately and coherently distributed over the surface of the Bloch sphere.
Figure 12: Selected examples visualizing single-qubit timestep embedding states. Our method allows the timestep embedding state to adaptively distribute itself on the surface of the Bloch sphere while maintaining temporal coherence. This demonstrates that our approach to encoding temporal information is expressive and flexible.
Compared to qubit encoding methods [58, 59, 60], our approach to timestep embedding, as inspired by Ref. [63], is more expressive and flexible. The qubit encoding method generates states that are equidistantly arranged in a ring on the Bloch sphere surface. This restricts subsequent learning models to only learning from such fixed patterns. In contrast, our approach allows the timestep embedding states to be adaptable, enhancing their multifunctionality. This could potentially decrease the difficulty of learning for subsequent models and improve the overall model’s performance.
VI-C Mixed State Generation
In Fig. 13, we evaluate the fidelity of four quantum generative models for generating mixed states. The results indicate that the fidelity of QuGAN and EQ-GAN decreases as the number of qubits increases, while QGDM and RE-QGDM still maintain high generation fidelity in large-scale tasks. Specifically, QuGAN exhibits a continuous decrease in fidelity with an increase in $N$, sharply falling from above 0.9 at $N=1$ to approximately 0.3 at $N=8$ for most fidelity points and the median. EQ-GAN also shows a significant decrease in fidelity but maintains a better median and data point distribution compared with QuGAN. These findings suggest that EQ-GAN outperforms QuGAN in generating mixed states, but both are difficult to handle the task of generating mixed states with multiple qubits. In contrast, QGDM and RE-QGDM consistently show a median fidelity above 0.99 across all scenarios. QGDM surpasses both QuGAN and EQ-GAN at $N=1$, with most fidelity points above 0.99. For $N=$2, 3, and 4, RE-QGDM has a slight advantage over QGDM, with RE-QGDM having more concentrated data points. Notably, at $N\geq 5$, RE-QGDM is significantly better than QuGAN and EQ-GAN, maintaining all fidelity scores above 0.9.
Figure 13: Mixed state generation results. The circles represent outliers, which are data points that significantly deviate from the rest of the data.
TABLE III: Comparison of mixed state generation performance between QuGAN, EQ-GAN, and our proposed methods.
Models $N=1$ $N=2$ $N=3$ $N=4$ $N=5$ $N=6$ $N=7$ $N=8$
QuGAN [13] 0.902$\pm$6e-2 0.694$\pm$1e-1 0.598$\pm$6e-2 0.477$\pm$9e-2 0.414$\pm$8e-2 0.439$\pm$9e-2 0.396$\pm$8e-2 0.324$\pm$6e-2
EQ-GAN [21] 0.972$\pm$3e-2 0.867$\pm$8e-2 0.741$\pm$1e-1 0.676$\pm$7e-2 0.602$\pm$6e-2 0.555$\pm$6e-2 0.463$\pm$1e-1 /
QGDM (Ours) 0.995$\pm$7e-3 0.999$\pm$1e-3 0.992$\pm$1e-2 0.992$\pm$7e-3 / / / /
RE-QGDM (Ours) / 0.999$\pm$6e-4 0.999$\pm$5e-4 0.996$\pm$2e-3 0.993$\pm$4e-3 0.990$\pm$1e-2 0.992$\pm$2e-2 0.993$\pm$2e-2
In Table. III, we present a detailed comparison of the fidelity among these four methods. QuGAN starts at a fidelity of 0.902$\pm$6e-2 at $N=1$, declining to 0.324$\pm$6e-2 at $N=8$. This trend highlights QuGAN’s limitations regarding scalability. EQ-GAN has a lower standard deviation in most cases (except for the case $N=3$), while generally displaying higher average fidelity compared to QuGAN. In contrast, QGDM and RE-QGDM maintain a fidelity of greater than 0.99 across all results, accompanied by a low standard deviation, indicating their superior scalability. Notably, our methods achieve fidelity of up to 0.999 in some cases, demonstrating their excellent generative capabilities. In particular, RE-QGDM achieves 0.992$\pm$1e-2 for $N=7$ and maintains 0.993$\pm$1e-2 for $N=8$, demonstrating the ability of our models to maintain high fidelity generation even on larger scale problems.
Our method demonstrated significant advantages in comparison with QuGAN and EQ-GAN. The results indicate that QuGAN and EQ-GAN perform poorly in the tasks of generating mixed states. The fidelity decreases as the number of qubits increases, revealing the limitations of these models in handling multi-qubit tasks. This limitation might be due to QGANs’ optimization strategies, which rely on the min-max saddle point game, which is generally less effective than using gradient descent for convex functions. It is well known that training classical GAN models is prone to issues such as model collapse, instability, and non-convergence. With an exponential increase in the dimensions of the density matrix, there may be higher demands on the training stage, such as the need for a more suitable learning rate. Given these conditions, the QGANs’ unstable training performance may significantly impact the final generation results. In contrast, the training process of (RE-) QGDM, which minimizes a convex function using gradient descent, typically shows more user-friendly and stable training performance. This might explain why our model outperforms QuGAN and EQ-GAN.
VII Models Training and Generation Curves
Figure 14: Metrics of training and generation process of QGDM.
Figure 15: Metrics of training and generation process of RE-QGDM when $N$ = 2, 3, 4 and 5.
Figure 16: Metrics of training and generation process of RE-QGDM when $N$ = 6, 7 and 8.
Fig. 14 displays the metrics during the training and generation of QGDM, which include the loss values and the fidelity of the generated quantum states. In contrast, Fig. 15 and Fig. 16 show the metrics of RE-QGDM.
Each scenario (number of qubits, pure or mixed state) includes two subfigures: the left subfigure illustrates the trend of training loss function value as the number of epochs increases, while the right subfigure shows how fidelity changes over timesteps. The red dashed line indicates the fidelity changes between the noisy quantum state and the target state during the diffusion process, as time steps increase (indicated by the red arrow from right to left). The solid blue line represents the trend of fidelity between the generated quantum state and the target state during the generation process (indicated by the blue arrow, from left to right). The shaded area in the figure represents the range of standard deviation.
The loss curves indicate that the loss functions of both QGDM and RE-QGDM converge to near zero as training epochs increase, and the number of epochs required to converge to zero increases with $N$, regardless of whether the target state is pure or mixed. Moreover, generating mixed states requires more iterations than pure states, especially as $N$ increases. This suggests that generating mixed states is more challenging than generating pure states.
As shown in Appendix B, the fidelity between any quantum state $\rho$ and the completely mixed state is $\frac{1}{d}\leq F(\rho,\mathbb{I}/d)\leq 1$. For pure states, the fidelity between the target state and the completely mixed state is $1/d$. Therefore, the standard deviation of all diffusion fidelity curves (red dashed lines) for pure states is zero. In the case of mixed states, the lower bound of the fidelity is $\frac{1}{d}$, and the variance decreases as $N$ increases. Therefore, at $t=T$ ($i.e.$, the far left of the horizontal axis), both the fidelity values and the standard deviation decrease with $N$.
In all scenarios, the fidelity curve during generation (blue solid line) increases as the timestep $t$ decreases, eventually approaching 1 with the smallest standard deviation. The overlap between the generation fidelity curve and the diffusion fidelity curve decreases as the number of qubits increases. As the number of qubits increases, the generation-fidelity curve diverges more from the diffusion-fidelity curve. Additionally, the generation fidelity curve (blue solid line) becomes steeper as $N$ increases.
VIII Conclusion
Inspired by classical diffusion models, this paper introduces the Quantum Generative Diffusion Model (QGDM), a novel quantum generative approach for generating a quantum state ensemble. The fundamental concept of QGDM is that any target quantum state can be converted into a completely mixed state through a non-unitary forward process. Following this, a trainable backward process can be employed to reconstruct the target state from the completely mixed state. In the forward process, a timestep-dependent depolarizing channel incrementally converts the input state into a maximally random state, known as a completely mixed state. The denoising process involves a parameterized function composed of a timestep embedding circuit and a denoising circuit, employing the concept of parameter sharing. Specifically, the timestep embedding circuit embeds temporal information into a quantum state, while the denoising circuit processes the polluted quantum state along with the timestep embedding state to produce a denoised state. To minimize the number of auxiliary qubits in QGDM, we introduce a novel resource-efficient version of QGDM (RE-QGDM). Numerical simulations, conducted up to 8 qubits, compare our models with QuGAN and EQ-GAN. The results indicate that both QGDM and RE-QGDM surpass QuGAN and EQ-GAN in generating both pure and mixed quantum states, highlighting their excellent performance and scalability.
Future work could explore the impact of different noise schedules on generation performance. Additionally, designing more efficient structures for the denoising process to reduce resource requirements, including the number of auxiliary qubits and the total number of quantum gates, is a valuable and worthwhile direction to pursue. Investigating the relationship between problem properties and the number of different time steps can provide important insights for further understanding QGDM. Theoretically analyzing and proving that QGDM has better convergence than previous QGAN models is an important direction for future exploration. This could provide theoretical support for a series of quantum versions of classic diffusion models and deepen our understanding of the advantages of these models.
Appendix A The hyperparameters of the proposed models
TABLE I: The hyperparameters of QGDM.
Parameters Pure state / Mixed state
$N=1$ $N=2$ $N=3$ $N=4$
$L_{\mathcal{T}}$ 5
$\lambda$ 0.02
$s$ 0.008
Batch size 16
$T$ 30
Epoch 200
Initial learning rate 0.3
Final learning rate 0.01
Learning rate decay steps 200
$L$ 1 1 / 2 2 / 3
TABLE II: The hyperparameters of RE-QGDM.
Parameters Pure state / Mixed state
$N=2$ $N=3$ $N=4$ $N=5$ $N=6$ $N=7$ $N=8$
$L_{\mathcal{T}}$ 5
$\lambda$ 0.02
$s$ 0.008
Batch size 16
$T$ 30 30 / 90
Epoch 200 200 / 500
Initial learning rate 0.3 0.5
Final learning rate 0.01 0.07
Learning rate decay steps 200 200 / 500
$L$ 1 1 / 2 1 / 3 1 / 4
Appendix B Derivation of fidelity boundary between a quantum state and completely mixed state
We start by considering the definition of fidelity $F$ between a quantum state $\rho$ and the completely mixed state $\frac{\mathbb{I}}{d}$, where $\mathbb{I}$ is the identity matrix and $d$ is the dimension of the Hilbert space. The fidelity is given by:
$F\left(\rho,\frac{\mathbb{I}}{d}\right)$ $=\left(\operatorname{tr}\sqrt{\sqrt{\rho}\frac{\mathbb{I}}{d}\sqrt{\rho}}\right)^{2}$ (A1)
$=\frac{1}{d}\left(\operatorname{tr}\sqrt{\rho}\right)^{2}$
$=\frac{1}{d}\left(\operatorname{tr}\sum_{k}\sqrt{\lambda_{k}}|k\rangle\langle k|\right)^{2}$
$=\frac{1}{d}\left(\sum_{i,k}\sqrt{\lambda_{k}}\langle i|k\rangle\langle k|i\rangle\right)^{2}$
$=\frac{1}{d}\left(\sum_{k}\sqrt{\lambda_{k}}\right)^{2}$
Three distinct cases need to be considered.
1. If $\rho$ is the completely mixed state, then fidelity reaches its maximum:
$F\left(\rho,\frac{\mathbb{I}}{d}\right)=1.$ (A2)
2. If $\rho$ is a mixed state that is not the completely mixed state, the eigenvalues $\{\lambda_{k}\}$ are not necessarily equal but are non-negative and sum to one. This results in:
$F\left(\rho,\frac{\mathbb{I}}{d}\right)$ $=\frac{1}{d}\left(\sum_{k}\sqrt{\lambda_{k}}\right)^{2}$ (A3)
$>\frac{1}{d}\sum_{k}\left(\sqrt{\lambda_{k}}\right)^{2}$
$=\frac{1}{d}.$
Thus, if $\rho$ is a mixed state that is not the completely mixed state, the fidelity holds $F\left(\rho,\frac{\mathbb{I}}{d}\right)>\frac{1}{d}$.
3. If $\rho$ is a pure state ($\rho=|\psi\rangle\langle\psi|$), where $\rho$ has a unique eigenvalue 1 and its corresponding eigenvector is $|\psi\rangle$. Then
$F\left(\rho,\frac{\mathbb{I}}{d}\right)=\frac{1}{d}$ (A4)
In summary, the fidelity boundary between any quantum state $\rho$ and the completely mixed state $\frac{\mathbb{I}}{d}$ is
$\frac{1}{d}\leq F\left(\rho,\frac{\mathbb{I}}{d}\right)\leq 1.$ (A5)